Get Help

Online: Request Help
Phone: (707) 826-HELP (4357)
Email: help@humboldt [dot] edu
In Person: Library 120 • Hours
Reset HSU Password
System Status

Services For

System Status Archive

System Issue start Who is affected What happened
CMS/Peoplesoft Jun 19 2013 - 18:00 All CFS Production (FCFSPRD) and CFS Data Warehouse application and database services users. During our regularly scheduled maintenance window the following items will be installed: PeopleSoft Fixes for 9.0 AP, AR Aging by Customer Report, Enterprise Fees, Instructional Equipment, Reporting Solution, and GAAP Reports.
Other Jun 18 2013 - 09:41 All Portal users HSU Portal server was temporarily not accessible for approximately 10 minutes. It is now back up and available.
Other Jun 17 2013 - 15:13 Sophos Home Users Sophos Home Users may have noticed in the past few days that there was a red "x" on the Sophos icon in their system trays indicating that Sophos could not contact the server. The system has been restarted and updating service is restored.
Other Jun 15 2013 - 13:00 All Nolij Users / OAK database server users Shutdown delays due to excessive Database server Memory Use.
Other Jun 15 2013 - 05:00 All Users Scheduled Maintenance
Off-Campus Internet Connection Jun 6 2013 - 09:03 All campus community ATT issue identified affecting multiple sites in Northern California. CENIC is working with ATT to resolve the issue.
Moodle Jun 8 2013 - 12:00 >> This Event Has Been Cancelled << Faculty & Students >> This Event Has Been Cancelled << Maintenance window.
CMS/Peoplesoft Jun 6 2013 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During an additional planned outage outside of our regularly scheduled maintenance window, the following items will be installed: DARS Interactive Changes and CAS Authentication. Special Notes related to CAS Authentication: 1. You may see the login page flash on your browser, before you are authenticated. 2. If you access HCM production without going through the Portal, you will now be redirected to the CAS login. 3. If you have a batch account, you will receive a separate login URL after the changes have been successfuly verified.
Domain Name Service Jun 5 2013 - 17:00 All Users Rescanning the RAID controller after adding a hot-spare drive unexpectedly led to a system state requiring a reboot. The server has been rebooted, everything looks well.
Other Jun 5 2013 - 14:30 Virtual Desktop Users (not UPD) Network reconfiguration
Other Jun 5 2013 - 09:24 Users of OBI and SDR system
CMS/Peoplesoft Jun 5 2013 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly schedule maintenance window the following item will be installed: Oracle 11g Upgrade. NOTE: The Absence Hint Optimization baseline change will go in on 6/6 at 10:00 a.m with no system outage.
Other May 30 2013 - 15:59 All campus users ITS is investigating, stay tuned. [Connectivity issues resolved-still investigating cause.]
CMS/Peoplesoft May 29 2013 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window the following items will be installed: MP 8.05, MP 8.06, MP 8.07, MP 8.08, MP 8.09, Finance Aid Self-Service Retrofit, Orientation Modification View Change, and Absence Search Improvement.
Campus Phone System May 21 2013 - 10:55 All campus phone users making off-campus local phone calls *** UPDATE - A power surge at the AT&T Central Office caused the problem, the circuit has been replaced and service is restored *** AT&T is looking into it. Please check back for updates.
Campus Phone System May 20 2013 - 08:10 All campus voicemail users *** UPDATE - System is working now, TNS is still investigating cause and monitoring system *** Unknown at this time. TNS is investigating. Check back for updates.
Campus Data Network May 15 2013 - 11:45 Any network user in the BSSB Forum **** UPS is now on house power. Service should be restored. Will report back when the UPS has new batteris. *** The batteries in the UPS have died. TNS is working to move power to house power until we can update the batteries in the ups.
Other May 13 2013 - 13:32 Campus Users Please ignore and delete. If any of your users fell for it, just have them reset their password. The Help Desk (x4357) can assist with this.
Off-Campus Internet Connection May 12 2013 - 19:47 All campus community *** UPDATE - Higher than normal average bandwidth usage coupled with a large burst of bandwidth at 7:47PM exceeded total available up-link capacity. This in turn resulted in BGP hold down timers expiring, causing a fail-over which didn't help the situation. We are looking at our options to prevent this from happening again in the future. **** It is still unknown what happened, however we continue to investigate possibilities and will post more information tomorrow, May 13th, 2013.
Google Apps - Mail May 10 2013 - 11:16 Campus Users Google disabled this very quickly so most likely no one has been affected, however if you DO have a user that entered their info before Google took it down, just have them reset their password. The Help Desk can assist (x4357).
Other May 3 2013 - 08:13 Less than 1% of Windows 7 Users Only Unknown at this time, so far 12 machines have been affected out of over 7000. TNS is working with the vender and affected machines to determine the cause.
Google Apps - Mail May 3 2013 - 10:51 Campus Users We are working on a global block.
Wireless Network Apr 23 2013 - 22:55 Trinidad Marine Lab, Brett Harte House, Wagner House Wireless & Wireless WMS service failed on the master controller.
Other Apr 21 2013 - 05:00 All Users Scheduled Maintenance Window - 3rd Sunday of the month Window has been extended until noon.
Campus Data Network Apr 21 2013 - 03:00 Any Key Advisor or ITC trying to access their web page (https://tns.humboldt.edu/keyadvisor or https://tns.humboldt.edu/itc) During this weekend's maintenance window, the server that holds these services is being virtualized. The system will be down for most of the day for this work.
Off-Campus Internet Connection Apr 16 2013 - 11:35 All campus Internet users trying to access off campus internet TNS is investigating. Check back for updates. ** TNS is still looking into issue with CENIC ** *** Issue was found to be caused by a Denial of Service attack ***
Other Apr 18 2013 - 13:55 Campus Users We are currently investigating. Stay tuned for updates. One of the redundant CAS nodes experienced an error, and has been isolated. CAS is working again. ITS will investigate the error logs, and re-enter the failed node into the mix during the Maintenance Window on Sunday.
Off-Campus Internet Connection Apr 17 2013 - 12:26 Trinidad Marine Lab. ONLY *** Service has been restored *** Service went down at 12:26 am. Service is still down. No ETR at this time has been given. Check back for updates.
CMS/Peoplesoft Apr 17 2013 - 18:00 Finance and Finance Data Warehouse users The following changes do not impact campus end-users other than the fact that the systems will be unavailable. CMS Central will install into CFS: Update to ProCard, Update to Data Security, Update to AR Lockbox with Campus FTP Directory.
Other Apr 16 2013 - 18:25 All users Single Sign On login system required cache reset.
feedback